OpenStudy (anonymous):

How does the division rule for exponents help in understanding why anything to the zero power is 1?

OpenStudy (dumbcow):

\[\frac{5}{5} = 1\] this shows that anything divided by itself is 0 \[\frac{x^2}{x^2} = x^{2-2} =x^0 =1\]

OpenStudy (dumbcow):

1*

OpenStudy (anonymous):

Ok, I think I get it now